OpenStudy (anonymous):

Challege anyone? A plumber and his assistant work together to replace the pipes in an old house. The plumber charges $40 an hour for his own labor and $15 an hour for his assistant's labor. The plumber works twice as long as his assistant on this job, and the labor charge on the final bill is $3325. How long did the plumber and his assistant work on this job?

OpenStudy (anonymous):

plumber's charge = $40 an hour assistant's charge = $15 an hour plumber's time working = 2 * that of his assistant Total charge: plumber's charge an hour * amount of hours worked + assistant's charge an hour * amount of hours worked Total charge: 40 * (2a) + 15 * a = 3325 80a + 15 a = 3325 95 a = 3325 19 a = 665 a = 665/19 = 35

OpenStudy (anonymous):

HOw much for plumbler then his assistant?

OpenStudy (anonymous):

a is the amount of time the assistant worked, hence the 2a as the time the plumber worked. So a = 35 = time assistant worked, time plumber worked = 70.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome!